Operations Report – February 5th

Crew 221 Operations Report 05Feb2022

SOL: 3

Name of person filing report: Luis D. Monge

Non-nominal systems: Generator

Notes on non-nominal systems: on-going issue

Generator: on (at 7pm)

Hours run: 11:00

From what time last night: 9:00pm

To what time this morning: 8:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 80%

Diesel Reading – 100%

Station Propane Reading – 61%

Ethanol Free Gasoline – gallons

Water (loft tank): 38 gallons

Water Meter: 148 units

Water (static tank) – 390 gals

Static to Loft Pump used – yes

Water in GreenHab – 252.7 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Bayo rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: (before EVA) : 129

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 35%

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: yes

Hours: 135

Beginning charge: 100%

Ending charge: 25%

Currently charging: yes

Notes on rovers: Opportunity still out of commission, brakes need fixing before being delivered to the Hab

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: Notification received on Router replacement, rescheduled to be performed tomorrow as per instructions.

Summary of suits and radios: Fan on suit 7 stoped working during EVA, it restarted after reset. It will be checked for losing conections before being used again.

Summary of Hab operations: full deep cleanig of the upper level of the hab carried out yesterday evening after a mouse was spoted next the couch.

Summary of GreenHab operations: operational, watering of the plants carried out as per procedure

Summary of ScienceDome operations: 2 incubators runing

Summary of RAM operations: not used

Summary of any observatory issues: dome was not properly closed after night observations, it was detected early in the morning reported and fixed.

Summary of health and safety issues: no issues

Questions, concerns and requests to Mission Support: a new mouse trap was requested and it was delivered early in the morning.

Copyright © The Mars Society. All rights reserved. | Main Site